The scores on a certain test are normally distributed with a mean score of 51 and a standard deviation of 2. What is the probability that a sample of 90 students will have a mean score of at least 51.2108
Answer by stanbon(75887)

The scores on a certain test are normally distributed with a mean score of 51 and a standard deviation of 2. What is the probability that a sample of 90 students will have a mean score of at least 51.2108
---
z(51.2108) = (51.2108-51)/[2/sqrt(90)] = 1
---
P(x >= 51.2108) = P(z >= 1) = 0.1587
=============================
Cheers,
Stan H.